OpenStudy (anonymous):

find the volume of a sphere with a radius of 2 feet. express the answer in terms of pie and then round the answer to the nearest whole number.

zepdrix (zepdrix):

Volume of a sphere is given by:\[\Large\bf\sf V\quad=\quad \frac{4}{3}\pi \color{royalblue}{r}^3\] So let's plug our given value into the expression:\[\Large\bf\sf V\quad=\quad \frac{4}{3}\pi \left(\color{royalblue}{2ft}\right)^3\]

zepdrix (zepdrix):

Then we'll need to simplify.

zepdrix (zepdrix):

Apply the cube to each term in the brackets (cube the units also).

zepdrix (zepdrix):

\[\Large\bf\sf V\quad=\quad \frac{4}{3}\pi\cdot2^3ft^3\]

zepdrix (zepdrix):

Do you understand how to simplify it from there?

OpenStudy (anonymous):

can you solve it